# FCS-HGNN: Flexible Multi-type Community Search in Heterogeneous Information Networks

### v2

@@ Abstract.
Community search is a personalized community discovery problem designed to identify densely connected subgraphs containing the query node. Recently, community search in heterogeneous information networks (HINs) has received considerable attention. Existing methods typically focus on modeling relationships in HINs through predefined meta-paths or user-specified relational constraints. However, metapath-based methods are primarily designed to identify single-type communities with nodes of the same type rather than multi-type communities involving nodes of different types. Constraint-based methods require users to have a good understanding of community patterns to define a suitable set of relational constraints, which increases the burden on users. In this paper, we propose FCS-HGNN, a novel method for flexibly identifying both single-type and multi-type communities in HINs. Specifically, FCS-HGNN extracts complementary information from different views and dynamically considers the contribution of each relation instead of treating them equally, thereby capturing more fine-grained heterogeneous information. Furthermore, to improve efficiency on large-scale graphs, we further propose LS-FCS-HGNN, which incorporates i) the neighbor sampling strategy to improve training efficiency, and ii) the depth-based heuristic search strategy to improve query efficiency. We conducted extensive experiments to demonstrate the superiority of our proposed methods over state-of-the-art methods, achieving average improvements of 14.3% and 11.1% on single-type and multi-type communities, respectively.
@@ 1. Introduction
Community search, a kind of query-dependent personalized community discovery problem designed to identify densely-connected subgraphs containing query nodes, has garnered considerable interest (Luo et al., 2020; Guo et al., 2021; Chen et al., 2023b; Cai et al., 2023; Chen and Gao, 2022; Miao et al., 2022; Wang et al., 2024; Chen et al., 2023a). Most existing studies primarily focus on homogeneous graphs associated with a single type of nodes and edges. However, many real-world data are naturally represented as heterogeneous information networks (HINs), where multiple types of entities and the relations between them are embodied by various types of nodes and edges, respectively. Compared with homogeneous graphs, HINs possess richer semantics and complex relationships, encapsulating more valuable community information.
@@
Recent advances in community search over HINs usually endeavor to model the relationships between nodes of different types through the predefined meta-paths (Fang et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) or user-specified relational constraints (Jian et al., 2020). However, as shown in Table 1, they suffer from four main limitations. (1) These metapath-based methods are designed to identify single-type communities with nodes of the same type, rather than multi-type communities involving nodes of different types. The single-type community does not fully exploit the valuable community information in HINs. For example, as shown in Figure 1, the single-type community composed of authors (blue circle) can only uncover collaboration information among authors. Instead, a multi-type community spanning authors, papers, and venues (red circle) may reflect the core of a research field. By analyzing such multi-type communities, we can not only understand collaboration patterns among authors but also delve deeper into research trends by papers and influential venues in this field, thereby more effectively supporting various downstream applications (Chakraborty et al., 2015; Li et al., 2018b; Gajewar and Das Sarma, 2012; Behrouz et al., 2022). (2) Constraint-based methods require users to have a good understanding of community patterns, in order to define a suitable set of relational constraints. Otherwise, without any guidance, they may fail to find any communities, which significantly increases the burden on end-users. (3) Both metapath-based and constraint-based methods suffer from the pattern inflexibility, which refers to the problem that they are based on predefined community patterns, such as $(k,\mathcal{P})$-core (Fang et al., 2020; Zhou et al., 2023; Liu et al., 2023), and $(k,\Psi)$-NMC (Jiang et al., 2022a), or the user-specified relational constraints. In real-world scenarios, the pattern of a community is flexible in nature, and the target communities may not conform to predefined community patterns or relational constraints, thus risking being overlooked. (4) Both metapath-based and constraint-based methods only focus on structural properties of HINs (Fang et al., 2020; Jian et al., 2020; Jiang et al., 2022a) or limited low-dimensional node features (Zhou et al., 2023; Liu et al., 2023). For example, (Zhou et al., 2023) only considers one-dimensional features (such as influence). In practical scenarios, HINs usually naturally endow nodes with high-dimensional features containing important information, which is a challenge for existing methods in HINs.
@@
e.g., (Zhou et al., 2023; Liu et al., 2023) ✓(Low Dim.) Predefined Predefined
@@
To address the above limitations, we propose FCS-HGNN, a query-driven heterogeneous graph neural network that can flexibly identify multi-type communities, as well as single-type communities in HINs. To the best of our knowledge, we are the first to propose a machine learning (ML) based community search algorithm tailored for heterogeneous information networks. To tackle pattern inflexibility, FCS-HGNN adaptively learns community patterns in a data-driven manner rather than relying on strict predefined community patterns or user-specified relational constraints. Specifically, we design two encoders: (1) Heterogeneous Encoder to deal with the heterogeneity and learn query-independent node features, and (2) Query Encoder to learn the local structural information around the query nodes. In each encoder, we introduce an edge semantic attention mechanism to dynamically consider the contribution of each relation when identifying diverse multi-type communities rather than treating them equally. It is worth noting that graph neural networks offer a unified way to combine the structural information and the arbitrary dimensional node features. To take advantage of complementary information from different encoders, FCS-HGNN integrates them through an attention mechanism after each layer and ultimately utilizes them in the final layer to infer the probability of each node belonging to the community. Finally, by considering the nodes’ probabilities belonging to the community, we employ a breadth-first search (BFS) based algorithm to obtain the target multi-type community.
Furthermore, to improve efficiency on large-scale graphs, we introduce the neighbor sampling algorithm and depth-based heuristic search strategy to further optimize our algorithm, named LS-FCS-HGNN (Large Scale FCS-HGNN). Through the neighbor sampling algorithm, in each iteration, LS-FCS-HGNN updates its parameters on a randomly sampling subgraph, rather than on the entire large-scale graph, which significantly reduces the computational resource requirements and improves the efficiency during the training phase. Furthermore, based on real-life observations, we find that the nodes within the same community typically exhibit more edges and shorter paths. In other words, the nodes closer to the query node are more likely to belong to the community. Therefore, we propose a depth-based heuristic search strategy to reduce searches for the nodes farther from the query node, thereby significantly improving query efficiency.
@@
$\bullet$ We further propose LS-FCS-HGNN, an advanced algorithm designed to improve the efficiency in large-scale graphs. LS-FCS-HGNN incorporates two strategies including i) the neighbor sampling algorithm to reduce the computational resource requirements and improve training efficiency, and ii) the depth-based heuristic strategy to improve the query efficiency, while preserving community effectiveness.
$\bullet$ We conduct extensive experiments on five real-world heterogeneous information networks to demonstrate the superiority of our proposed methods over state-of-the-art methods. Our proposed FCS-HGNN and LS-FCS-HGNN significantly enhance the effectiveness of multi-type community search and improve the training and query efficiency.
Roadmap. The rest of the paper is organized as follows. Section 2 discusses related work. Section 3 provides some preliminaries. Section 4 details our proposed FCS-HGNN. Section 5 elaborates on our advanced LS-FCS-HGNN. Section 6 conducts comprehensive experiments to demonstrate the effectiveness of our proposed methods. Finally, section 7 concludes the paper.
@@ 2. RELATED WORK
Community Search. Over the past few decades, there has been extensive research on community search, with the primary objective of identifying a densely connected subgraph that encompasses the query nodes. To measure the structural cohesiveness of communities, various community patterns have been proposed, such as $k$-core (Seidman, 1983; Batagelj and Zaversnik, 2003), $k$-truss (Cohen, 2008; Huang et al., 2014), $k$-clique (Tsourakakis, 2015), and $k$-edge connected component (ECC) (Chang et al., 2015; Hu et al., 2016). However, the patterns of communities in real-world scenarios are flexible in nature. If real-world communities do not follow any of the above patterns, these models will fail to discover the true communities. Therefore, machine learning (ML)-based community search methods have been proposed (Gao et al., 2021; Jiang et al., 2022b; Li et al., 2023; Fang et al., 2023; Behrouz and Hashemi, 2022; Hashemi et al., 2023) to address the limitation of pattern inflexibility. However, most of these works primarily focus on homogeneous graphs where nodes belong to the same type.
@@
Recently, an increasing number of researchers have shown interest in community search over heterogeneous information networks (HINs) (Fang et al., 2021, 2020; Jian et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023), due to the prevalence and significance of HINs in real-world scenarios. These studies usually model the relationships between nodes in HINs through predefined meta-paths or user-specified relational constraints. For example, (Fang et al., 2020) proposes to employ predefined meta-paths to delineate the connectivity among nodes of different types in HINs, and extend the $k$-core to $(k,\mathcal{P})$-core to measure the cohesiveness of a community, where $\mathcal{P}$ is the predefined meta-path. Metapath-based methods aim to find single-type communities consisting of nodes with the same type, rather than discovering multi-type communities involving nodes of different types. (Jian et al., 2020) proposes a relational community model, which defines communities based on a set of user-specified relational constraints. It requires users to have a good understanding of community patterns to define a suitable set of constraints, which significantly increases the burden on users. If the constraints are not specified correctly, it may fail to find any communities. In contrast, our FCS-HGNN alleviates pattern inflexibility by adaptively learning community patterns in a data-driven manner, thereby alleviating the burden on end-users.
@@
Heterogeneous Graph Representation Learning. A large amount of work (Wang et al., 2022) has been devoted to address the problem of learning node representations from HINs, including heterogeneous graph embedding (Dong et al., 2017; Fu et al., 2017) and heterogeneous graph neural networks (Wang et al., 2019; Zhang et al., 2019; Lv et al., 2021; Yang et al., 2023). Heterogeneous graph embedding primarily focuses on preserving structure information based on meta-paths. For example, metapath2vec (Dong et al., 2017) devised a metapath-based random walk and utilized skip-gram (Mikolov et al., 2013) to learn node representations. Furthermore, according to the way of dealing with heterogeneity, heterogeneous graph neural networks can be divided into metapath-based methods and metapath-free methods. The design and selection of meta-paths still pose challenges in practice, which may require sufficient domain knowledge. When identifying multi-type communities, predefined meta-paths may lack flexibility, as different meta-paths are required for communities involving different user-specified node types. In contrast, our proposed FCS-HGNN adopts a metapath-free manner (Heterogeneous Encoder) to overcome heterogeneity and employs the Query Encoder to learn local structural information around query nodes, which sufficiently facilitates the learning of node representations in HINs.
@@ 3. Preliminaries
In this section, we first define heterogeneous information networks (HINs) and the multi-type community. Then, we formalize the multi-type community search problem in HINs.
Definition 0 (Heterogeneous Information Network).
A heterogeneous information network is a directed graph $\mathcal{H}=(V,E,\psi,\phi)$ with a node type mapping function $\psi:V\rightarrow\mathcal{A}$ and an edge type mapping function $\phi:E\rightarrow\mathcal{R}$, where $V$ and $E$ represent the sets of nodes and edges respectively, each node $v\in V$ belongs to a node type $\psi(v)\in\mathcal{A}$, each edge $e\in E$ belongs to an edge type $\phi(e)\in\mathcal{R}$, and $|\mathcal{A}|+|\mathcal{R}|>2$.
Due to its ability to delineate multiple types of nodes and edges as well as complex relationships, HINs have become ubiquitous in real-world scenarios, such as bibliographic networks (Sun et al., 2011; Shi et al., 2012a), IMDB movie networks (Shi et al., 2012b; Yu et al., 2013), and knowledge graphs (Hogan et al., 2021). Compared to homogeneous graphs, HINs encapsulate richer community information. Most existing methods primarily focus on single-type communities by meta-paths, which may not fully explore valuable community information in HINs. Therefore, we propose the multi-type community involving nodes of different types as follows:
@@
Given a HIN $\mathcal{H}=(V,E,\psi,\phi)$ and the set of node types $\mathcal{S}_{\mathcal{A}}\subseteq\mathcal{A}$, a multi-type community is defined as a subgraph $\mathcal{C}_{q}=(V_{q},E_{q},\psi,\phi)$ of $\mathcal{H}$ that satisfies the following criteria:
  1. (1)
    Each node $v\in V_{q}$ satisfies $\psi(v)\in\mathcal{S}_{\mathcal{A}}$;
  2. (2)
    $\mathcal{C}_{q}$ is connected and has a cohesive structure.
According to definition 2, the single-type community is a special case (when $|\mathcal{S}_{\mathcal{A}}|=1$) of the multi-type community. Metapath-based methods can identify single-type communities by predefined meta-paths, while constraint-based methods can identify multi-type communities by defining a set of relational constraints, as shown in Table 1. In contrast, our proposed FCS-HGNN can flexibly identify both single-type communities and multi-type communities without relying on strictly predefined meta-paths or relational constraints. Therefore, we formalize the problem statement for multi-type community search in HINs.
@@
Given a HIN $\mathcal{H}=(V,E,\psi,\phi)$, a query node $v_{q}\in V$, and the set of node types $\mathcal{S}_{\mathcal{A}}$, multi-type community search aims to identify the query-dependent community $\mathcal{C}_{q}=(V_{q},E_{q},\psi,\phi)$ satisfying Multi-type Community (refer to Definition 2) and $v_{q}\in V_{q}$.
Based on the above problem definition, we formalized the machine learning-based training manner as follows: given the set of node types $\mathcal{S}_{\mathcal{A}}$ and a set of query tasks with labeled communities, denoted as $\mathcal{D}=\{(q_{1},pos(q_{1}),neg(q_{1})),(q_{2},pos(q_{2}),neg(q_{2})),\cdots\}$, where $q$ represents a query node, $pos(q)$ is the set of positive nodes in the target community, $neg(q)$ is the set of negative nodes not in the target community. Our training objective is to learn a function $\mathcal{F}(\mathcal{H},q,\mathcal{S}_{\mathcal{A}})$ that yields the probability of each node belonging to the target community. Then, we can employ BFS to find the target community with the highest probability. It is worth noting that $neg(q)$ can be randomly sampled from the nodes except for the query node and $pos(q)$, without incurring additional annotation costs.

### v3

@@ Abstract.
Community search is a personalized community discovery problem designed to identify densely connected subgraphs containing the query node. Recently, community search in heterogeneous information networks (HINs) has received considerable attention. Existing methods typically focus on modeling relationships in HINs through predefined meta-paths or user-specified relational constraints. However, metapath-based methods are primarily designed to identify single-type communities with nodes of the same type rather than multi-type communities involving nodes of different types. Constraint-based methods require users to have a good understanding of community patterns to define a suitable set of relational constraints, which increases the burden on users. In this paper, we propose FCS-HGNN, a novel method for flexibly identifying both single-type and multi-type communities in HINs. Specifically, FCS-HGNN extracts complementary information from different views and dynamically considers the contribution of each relation instead of treating them equally, thereby capturing more fine-grained heterogeneous information. Furthermore, to improve efficiency on large-scale graphs, we further propose LS-FCS-HGNN, which incorporates i) the neighbor sampling strategy to improve training efficiency, and ii) the depth-based heuristic search strategy to improve query efficiency. We conducted extensive experiments to demonstrate the superiority of our proposed methods over state-of-the-art methods, achieving average improvements of 14.3% and 11.1% on single-type and multi-type communities, respectively. [^1] The code is available at https://github.com/Chen-GX/FCS-HGNN
Community Search, Multi-type Community
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679696isbn: 979-8-4007-0436-9/24/10ccs: Computing methodologies Machine learning
@@ 1. Introduction
Community search, a kind of query-dependent personalized community discovery problem designed to identify densely-connected subgraphs containing query nodes, has garnered considerable interest (Luo et al., 2020; Guo et al., 2021; Chen et al., 2023d; Cai et al., 2023; Chen and Gao, 2022; Miao et al., 2022; Wang et al., 2024; Chen et al., 2023a). Most existing studies primarily focus on homogeneous graphs associated with a single type of nodes and edges. However, many real-world data are naturally represented as heterogeneous information networks (HINs), where multiple types of entities and the relations between them are embodied by various types of nodes and edges, respectively. Compared with homogeneous graphs, HINs possess richer semantics and complex relationships, encapsulating more valuable community information.
@@
Recent advances in community search over HINs usually endeavor to model the relationships between nodes of different types through the predefined meta-paths (Fang et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) or user-specified relational constraints (Jian et al., 2020). However, as shown in Table 1, they suffer from four main limitations. (1) These metapath-based methods are primarily designed to identify single-type communities with nodes of the same type, rather than multi-type communities involving nodes of different types (Jian et al., 2020). The single-type community does not fully exploit the valuable community information in HINs. For example, as shown in Figure 1, the single-type community composed of authors (blue circle) can only uncover collaboration information among authors. Instead, a multi-type community spanning authors, papers, and venues (red circle) may reflect the core of a research field. By analyzing such multi-type communities, we can not only understand collaboration patterns among authors but also delve deeper into research trends by papers and influential venues in this field, thereby more effectively supporting various downstream applications (Chakraborty et al., 2015; Li et al., 2018b; Gajewar and Das Sarma, 2012; Behrouz et al., 2022; Chen et al., 2023b, 2024b, 2024a). (2) Constraint-based methods require users to have a good understanding of community patterns, in order to define a suitable set of relational constraints. Otherwise, without any guidance, they may fail to find any communities, which significantly increases the burden on end-users. (3) Both metapath-based and constraint-based methods suffer from the pattern inflexibility, which refers to the problem that they are based on predefined community patterns, such as $(k,\mathcal{P})$-core (Fang et al., 2020; Zhou et al., 2023; Liu et al., 2023), and $(k,\Psi)$-NMC (Jiang et al., 2022a), or the user-specified relational constraints. In real-world scenarios, the pattern of a community is flexible in nature, and the target communities may not conform to predefined community patterns or relational constraints, thus risking being overlooked. (4) Both metapath-based and constraint-based methods only focus on structural properties of HINs (Fang et al., 2020; Jian et al., 2020; Jiang et al., 2022a) or limited low-dimensional node features (Zhou et al., 2023; Liu et al., 2023). For example, (Zhou et al., 2023) only considers one-dimensional features (such as influence). In practical scenarios, HINs usually naturally endow nodes with high-dimensional features containing important information, which is a challenge for existing methods in HINs.
@@
e.g., (Zhou et al., 2023; Liu et al., 2023) (Low Dim.) Predefined Predefined
@@
To address the above limitations, we propose FCS-HGNN, a query-driven heterogeneous graph neural network that can flexibly identify multi-type communities, as well as single-type communities in HINs. To tackle pattern inflexibility, FCS-HGNN adaptively learns community patterns in a data-driven manner rather than relying on strict predefined community patterns or user-specified relational constraints. Specifically, we design two encoders: (1) Heterogeneous Encoder to deal with the heterogeneity and learn query-independent node features, and (2) Query Encoder to learn the local structural information around the query nodes. In each encoder, we introduce an edge semantic attention mechanism to dynamically consider the contribution of each relation when identifying diverse multi-type communities rather than treating them equally. It is worth noting that graph neural networks offer a unified way to combine the structural information and the arbitrary dimensional node features. To take advantage of complementary information from different encoders, FCS-HGNN integrates them through an attention mechanism after each layer and ultimately utilizes them in the final layer to infer the probability of each node belonging to the community. Finally, by considering the nodes’ probabilities, we employ a breadth-first search (BFS) algorithm to obtain the target multi-type community.
Furthermore, to improve efficiency on large-scale graphs, we introduce the neighbor sampling algorithm and depth-based heuristic search strategy to further optimize our algorithm, named LS-FCS-HGNN (Large Scale FCS-HGNN). Through the neighbor sampling algorithm, in each iteration, LS-FCS-HGNN updates its parameters on a randomly sampling subgraph, rather than on the entire large-scale graph, which significantly reduces the computational overhead and improves training efficiency. Furthermore, based on real-life observations, we find that the nodes within the same community typically exhibit more edges and shorter paths. In other words, the nodes closer to the query node are more likely to belong to the community. Therefore, we propose a depth-based heuristic search strategy to reduce searches for the nodes farther from the query node, thereby significantly improving query efficiency.
@@
$\bullet$ We further propose LS-FCS-HGNN, an advanced algorithm designed to improve the efficiency in large-scale graphs. LS-FCS-HGNN incorporates two strategies including i) the neighbor sampling algorithm to reduce computational overhead and improve training efficiency, and ii) depth-based heuristic strategy to improve the query efficiency, while preserving community effectiveness.
$\bullet$ We conduct extensive experiments on five real-world heterogeneous information networks to demonstrate the superiority of our proposed methods over state-of-the-art methods. Our proposed LS-FCS-HGNN not only significantly enhance the effectiveness and flexibility of multi-typed community search but also obviously improve both training and query efficiency.
@@ 2. RELATED WORK
Community Search in HINs. Recently, an increasing number of researchers have shown interest in community search over HINs (Fang et al., 2021, 2020; Jian et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023; Chen et al., 2023c), due to the prevalence and significance of HINs in real-world scenarios. These studies usually model the relationships between nodes in HINs through predefined meta-paths or user-specified relational constraints. For example, (Fang et al., 2020) proposes to employ predefined meta-paths to delineate the connectivity among nodes of different types in HINs, and extend the $k$-core to $(k,\mathcal{P})$-core to measure the cohesiveness of a community, where $\mathcal{P}$ is the predefined meta-path. Metapath-based methods aim to find single-type communities consisting of nodes with the same type, rather than discovering multi-type communities involving nodes of different types. (Jian et al., 2020) proposes a relational community model, which defines communities based on a set of user-specified relational constraints. It requires users to have a good understanding of community patterns to define a suitable set of constraints, which significantly increases the burden on users. If the constraints are not specified correctly, it may fail to find any communities. In contrast, our FCS-HGNN alleviates pattern inflexibility by adaptively learning community patterns in a data-driven manner, thereby alleviating the burden on end-users.
@@
Heterogeneous Graph Representation Learning. A large amount of work (Wang et al., 2022) has been devoted to address the problem of learning node representations from HINs, including heterogeneous graph embedding (Dong et al., 2017; Fu et al., 2017) and heterogeneous graph neural networks (Wang et al., 2019; Zhang et al., 2019; Lv et al., 2021; Yang et al., 2023). Heterogeneous graph embedding primarily focuses on preserving structure information based on meta-paths. For example, metapath2vec (Dong et al., 2017) devised a metapath-based random walk and utilized skip-gram (Mikolov et al., 2013) to learn node representations. Heterogeneous graph neural networks can be divided into metapath-based methods and metapath-free methods. The design and selection of meta-paths still pose challenges in practice, which may require sufficient expert knowledge. In contrast, our proposed FCS-HGNN adopts a metapath-free manner (Heterogeneous Encoder) to deal with heterogeneity and employs the Query Encoder to learn local structural information around query nodes, which sufficiently facilitates the learning of community information in HINs.
@@ 3. Preliminaries
Definition 0 (Heterogeneous Information Network, HIN).
A HIN is a directed graph $\mathcal{H}=(V,E,\psi,\phi)$ with a node type mapping function $\psi:V\rightarrow\mathcal{A}$ and an edge type mapping function $\phi:E\rightarrow\mathcal{R}$, where $V$ and $E$ represent the sets of nodes and edges respectively, each node $v\in V$ belongs to a node type $\psi(v)\in\mathcal{A}$, each edge $e\in E$ belongs to an edge type $\phi(e)\in\mathcal{R}$, and $|\mathcal{A}|+|\mathcal{R}|>2$.
@@
Given a HIN $\mathcal{H}=(V,E,\psi,\phi)$ and the set of node types $\mathcal{S}_{\mathcal{A}}\subseteq\mathcal{A}$, a multi-type community is defined as a subgraph $\mathcal{C}_{q}=(V_{q},E_{q},\psi,\phi)$ of $\mathcal{H}$ that satisfies the following criteria: (1) Each node $v\in V_{q}$ satisfies $\psi(v)\in\mathcal{S}_{\mathcal{A}}$; (2) $\mathcal{C}_{q}$ is connected and has a cohesive structure.
@@
Given a HIN $\mathcal{H}=(V,E,\psi,\phi)$, a query node $v_{q}\in V$, and the set of node types $\mathcal{S}_{\mathcal{A}}$, multi-type community search problem aims to identify the query-dependent community $\mathcal{C}_{q}=(V_{q},E_{q},\psi,\phi)$ satisfying Multi-type Community (refer to Definition 2) and $v_{q}\in V_{q}$.
Based on the above problem definition, we formalized the ML-based training manner as follows: given the set of node types $\mathcal{S}_{\mathcal{A}}$ and a set of query tasks $\mathcal{D}=\{(q_{1},pos(q_{1}),neg(q_{1})),(q_{2},pos(q_{2}),$ $neg(q_{2})),\cdots\}$, where $q$ represents the query node, $pos(q)$ and $neg(q)$ respectively represent the subsets of nodes that are in or not in the target community. Our training objective is to learn a function $\mathcal{F}(\mathcal{H},q,\mathcal{S}_{\mathcal{A}})$ that yields the probability of each node belonging to the target community. Then, we employ search algorithm to find the target community with the highest probability.